5-x=7+2x-4
how do you write this out to find x?

Answer by stanbon(75887)

5-x=7+2x-4
how do you write this out to find x?
--------------
Simplify the right side as follows:
5-x = 3+2x
Add x to both sides to get:
5 = 3+3x
Subtract 3 from both sides to get:
2=3x
Divide both sides by 3 and switch the sides to get:
x=2/3
-------------
Cheers,